2 t=0 4 13 2 12V 2H IL

Determine the value of the current in the inductor for the time = 150 ms
Answer #1

1 Using haplace prausform in tu cka 2 4 VA 1245 25 nodal Applying Analysis, at Node A 12 VA VA-O + S VA Ć 25+3 2 VA VA VA 6 2(245+ 36) VA [ 165 + 36] ulor VA= 6 (245+36) S (16st 36) find current across indutor. :- 12 VA 25+3 17 6 (245+36) (165+36) (2(6) x (3) (S) X (45+9) 189 18 (S) (45+9) (5) Xorx (8 + 9/4) 2 1 9 2 X () X (5+914) A B -+ s+ 914 Als=0 9 A = 2 1 (2jx (otaly)